[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri May 26, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri May 26 18:55:12 UTC 2017


HSU Shift Log for Fri May 26 2017 14:55:09 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: FSU/Halperin guidance is having some issues with  
UKMET data, which is affecting UKMET and consensus probabilities.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170526/2d2427bf/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list